Hi,
If you are using a table that considers the left region,
carry out these steps:
1. Convert P(Z > a) to the P(Z < a) form, e.g.,
P(Z > a) = 0.0838 becomes P(Z < a) = 1 - 0.0838 =
0.9162.
2. Search for the probability value in the table, e.g.,
0.9162.
3. Identify the row and column values corresponding to
the probability value, e.g., 1.3 is the row value and 8 is the
column value corresponding to 0.9162.
4. a is made up of these two values, e.g., a = 1.38.
If you have a graphic calculator, it'll be much easier to
find the value of a. Thanks.
Cheers,
Wen Shih